# spacing=nonfrench

### v2

@@ Abstract
The adoption of large cloud-based models for inference in privacy-sensitive domains, such as homeless care systems and medical imaging, raises concerns about end-user data privacy. A common solution is adding locally differentially private (LDP) noise to queries before transmission, but this often reduces utility. LDPKiT, which stands for Local Differentially-Private and Utility-Preserving Inference via Knowledge Transfer, addresses the concern by generating a privacy-preserving inference dataset aligned with the private data distribution. This dataset is used to train a reliable local model for inference on sensitive inputs. LDPKiT employs a two-layer noise injection framework that leverages LDP and its post-processing property to create a privacy-protected inference dataset. The first layer ensures privacy, while the second layer helps to recover utility by creating a sufficiently large dataset for subsequent local model extraction using noisy labels returned from a cloud model on privacy-protected noisy inputs. Our experiments on Fashion-MNIST, SVHN and PathMNIST medical datasets demonstrate that LDPKiT effectively improves utility while preserving privacy. Moreover, the benefits of using LDPKiT increase at higher, more privacy-protective noise levels. For instance, on SVHN, LDPKiT achieves similar inference accuracy with $\epsilon=1.25$ as it does with $\epsilon=2.0$, providing stronger privacy guarantees with less than a 2% drop in accuracy. Furthermore, we perform extensive sensitivity analyses to evaluate the impact of dataset sizes on LDPKiT’s effectiveness and systematically analyze the latent space representations to offer a theoretical explanation for its accuracy improvements. Lastly, we qualitatively and quantitatively demonstrate that the type of knowledge distillation performed by LDPKiT is ethical and fundamentally distinct from adversarial model extraction attacks.
1 Introduction
Many high-performance machine learning (ML) applications are trained and hosted by online services, which have better access to large datasets and the computation necessary to train large models. However, users of such models may be concerned about the leakage of sensitive information in the queries they make to such cloud providers during inference, who could monitor or inspect their queries and use the information for purposes without their knowledge. Instances of providers misusing users’ data abound. In 2016, for example, Yahoo secretly complied with the US government’s digital communication surveillance and used its custom spam and child pornography detection system to monitor the user’s emails [1, 2, 3]. In 2022, the Amazon Ring Doorbell was found to have disclosed users’ video and audio footage to the police without authorization [4]. Even the recent large language models have incurred data leakage [5, 6], leading to many organizations banning their use. Even with a trusted model provider, a compromised cloud platform could enable an adversary to infer the user’s queries via side-channels [7].
Such concerns become even more acute in the domain of applications involving privacy-sensitive data, such as medical imaging [8], disease diagnosis [9], financial fraud detection [10], and systems supporting the vulnerable populations such as the homeless [11], which increasingly rely on machine learning automated decision making. Further, such domains may be challenged by a scarcity of publicly available training data, making it difficult for users to obtain and train reliable models of their own. For instance, when a person facing a crisis enquires for homeless care support, their personal status and needs information may be used to determine their eligibility for shelter, but their data should not be used for surveillance or targeting of vulnerable individuals. Similarly, in a medical imaging system, a patient’s sensitive data may be used for diagnostic purposes, but they might be concerned about their data being shared with insurance providers or potential employers.
@@ 1 Introduction
Various methods for protecting privacy during inference have been proposed. For example, homomorphic encryption [12, 13] and hardware-enforced trusted execution environments [14] can help to protect users’ inference queries against a compromised cloud platform. However, they cannot protect against a malicious or faulty model that may leak information about the inputs it is performing inference on.
An alternative approach that obtains strong privacy guarantees is to limit the amount of information that is sent to the model in a query. One way to limit the amount of information is to use a local differential privacy (LDP) based inference mechanism, which bounds the amount of information leakage by adding random noise to the query inputs before transmitting them for inference [15]. However, to obtain a reasonable guarantee of privacy, LDP requires a large amount of noise to be added to each input, leading to loss of utility [16].
In an ideal case, users could just train and use their own model locally, or a model provider could share the weights with a user enabling them to use the model locally. Such a model frees users from having to send any sensitive data to a model provider. However, in many specialized scenarios, it may be challenging for users to obtain the large, labelled training sets that are required to train highly accurate models. Similarly, in addition to economic motivations, model providers with models trained on private data may be reluctant to share their models with users out of concern about leaking the sensitive data in the model itself [17].
In this work, we propose, LDPKiT (Local Differentially-Private and Utility-Preserving Inference via Knowledge Transfer), which uses LDP to enable users to obtain a local model through knowledge transfer (i.e., model extraction). A key challenge with model extraction is that the user requires a set of in-distribution data with which to query the model. We seek to address scenarios where the availability of public data in the same distribution as the user’s private data is scarce, thus forcing the user to use their own private data for extraction. Naïvely adding LDP to the data before querying introduces significant noise into the labels returned by the model, degrading the model extraction. LDPKiT overcomes this challenge by adding noise to each private input, and then superimposing pairs of the noised inputs with each other before querying them. This has the benefit of creating more in-distribution inputs, enabling efficient knowledge transfer from the model provider.
@@
Model extraction can be a controversial technique. In the extreme case, malicious model extraction (i.e., model theft), where an adversary extracts a model provider’s commercial model and then resells that model can cause economic harm to the model provider and disincentivize model providers from training or providing machine learning services. We naturally do not advocate such uses of model extraction. In contrast, our approach differs from this model theft in several important ways: 1) LDPKiT’s extracted model’s performance does not match the provider’s model due to the addition of LDP noise. As a result, the model extraction is not for competitive purposes, and would not violate the terms of service of most commercial model providers [18]—the main objective of LDPKiT’s model extraction is to protect user privacy. 2) Users still perform queries to the model provider, and presumably compensate the model provider for these queries. As preventing model extraction, especially for small models, is inherently challenging, it is often accounted for in query pricing structures [19, 20, 21]. We further discuss our ethical use of model extraction principles in detail in Section 6.1.
@@
RQ1. Does LDPKiT recover the utility impacted by LDP noise? (Section 4.2)
RQ2. Why superimposition works in LDPKiT? (Section 4.3)
@@
RQ3. How do the size of the private dataset and number of queries impact LDPKiT? (Section 4.4)
Contributions.
@@
LDPKiT generates inference datasets that are both privacy-protective and close to the distribution of sensitive datasets, allowing users to accurately annotate their private data with mitigated privacy concerns. We evaluate LDPKiT across diverse applications and multiple model architectures, systematically examining its performance and sensitivity to dataset sizes. Additionally, we conduct in-depth analyses of latent space representations to uncover the underlying factors contributing to LDPKiT’s effectiveness.
2 Overview
@@
Figure 1: Comparison of accuracies: Training a local model (ResNet-18) with out-of-distribution (OOD) public data samples versus querying the cloud service model with $\epsilon$-LDP protected private data samples.
2.1 Target Scenario
Our target scenario is the use of privacy-sensitive ML cloud services. As discussed in Section 1, the ideal, but unrealistic, solution would be granting users access to cloud-hosted models for local deployment. Instead, users may opt to train their own models through knowledge transfer techniques. We target privacy-sensitive applications whose in-distribution public data is not readily available. As an illustration, we train two instances of ResNet-18 model with one of the two datasets, SVHN and Fashion-MNIST, and subsequently run inference queries on each model using data from the dataset on which it was not trained. As shown in Figure 1, the model accuracy drops to the level of random guessing (i.e., around 10% for 10-class datasets) when it is extracted using out-of-distribution (OOD) data samples (e.g., SVHN samples are considered OOD for the Fashion-MNIST task, and vice versa). Instead of extracting a model from OOD samples and recording its accuracies, we focus on querying a model (e.g., ResNet-152 in this experiment) trained on in-distribution data using queries derived from the same datasets but modified with LDP under highly privacy-protective $\epsilon$ settings. Even without any utility recovery techniques, the recorded accuracies based on the returned labels remain around 20% on average. This observation motivates the use of $\epsilon$-LDP as a viable starting point in our privacy-preserving framework, which we further enhance with techniques to recover the accuracy loss from the LDP noise. Further details are provided in Section 3.
Our design goals are to protect sensitive user queries when using ML cloud services during inference time and to recover some accuracy loss due to privacy-protective LDP noise added to the queries. We assume the cloud model provider is honest but curious. It honestly answers the user’s queries but may record both the queries and their results to infer information about the user. To evaluate the framework in a more challenging scenario, we also restrict access to the cloud model by assuming that the cloud model returns hard labels only.
2.2 Preliminaries and General Setup
Throughout this paper, we use ${{\mathcal{M}_{\rm R}}}$ to denote the remote, cloud-hosted model and ${{\mathcal{M}_{\rm L}}}$ to denote the local model hosted by the user. We use ${\mathcal{D}_{\rm priv}}$ to denote the sensitive dataset the user owns and wishes to protect. Note that ${\mathcal{D}_{\rm priv}}$ can be a predefined or dynamically generated set of points. We denote the number of private data points (i.e., size of ${\mathcal{D}_{\rm priv}}$) as $|{\mathcal{D}_{\rm priv}}|$. ${\mathcal{D}_{\rm protected}}$ refers to the privacy-protected ${\mathcal{D}_{\rm priv}}$ with a single layer of $\epsilon$-LDP noise applied. It has the same size as ${\mathcal{D}_{\rm priv}}$. The complete set of candidates derived from post-processing ${\mathcal{D}_{\rm protected}}$ with LDPKiT is represented as ${\mathcal{D}_{\rm cand}}$, with its size $|{\mathcal{D}_{\rm cand}}|$ calculated as $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$. Detailed explanations regarding this process are provided in Section 3. Subsequently, the user can construct a ${\mathcal{D}_{\rm infer}}$ by selecting candidate data points from ${\mathcal{D}_{\rm cand}}$ to query ${\mathcal{M}_{\rm R}}$ for labels. These labels are used to train an ${\mathcal{M}_{\rm L}}$ to annotate ${\mathcal{D}_{\rm priv}}$. The size of ${\mathcal{D}_{\rm infer}}$ ($|{\mathcal{D}_{\rm infer}}|$), representing the number of queries made, is determined by the user, with a maximum possible size of $|{\mathcal{D}_{\rm cand}}|=|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$. We use SIDP to denote the process of the standard inference scheme with LDP where a single layer of noise is added to ${\mathcal{D}_{\rm priv}}$ to form ${\mathcal{D}_{\rm protected}}$. The user then queries ${\mathcal{M}_{\rm R}}$ using ${\mathcal{D}_{\rm protected}}$ to obtain labels, which may include some errors due to the privacy-protective noise in the queries. In the SIDP method, the user straightforwardly accepts the potentially erroneous labels provided by ${\mathcal{M}_{\rm R}}$ without further post-processing. For instance, the LDP approach with results illustrated in the two rightmost scenarios of Figrue 1 represents an SIDP usage. Although SIDP is promising in privacy preservation, it degrades utility significantly. Instead of accepting the errors, we propose two ways to post-process ${\mathcal{D}_{\rm protected}}$ used in SIDP for our accurate privacy-preserving inference mechanism. We denote them as LDPKiT-Rand and LDPKiT-Sup. The details are discussed in Section 3.
2.3 Privacy Guarantee
@@
LDPKiT ensures the privacy of ${\mathcal{D}_{\rm priv}}$ during inference by applying LDP noise to each data sample. This section discusses the key concepts and methodologies essential for designing effective privacy-preserving noise mechanisms.
Definition 2.1.
@@
A lower $\epsilon$ value indicates a tighter bound of the equation and a stronger privacy guarantee. As for the choice of random noise, we apply Laplacian noise to satisfy the $\epsilon$-LDP privacy guarantee and the definition is as follows.
Definition 2.2.
@@
For the design of the candidate dataset used for inference, ${\mathcal{D}_{\rm cand}}$, we also leverage LDP’s post-processing property to facilitate utility recovery. If an algorithm safeguards an individual’s privacy, a data analyst cannot compromise privacy—whether under the formal definition or in any intuitive sense—merely by reflecting on the algorithm’s output [23]. Post-processing is commonly used in DP schemes to improve the interpretability or accuracy of differentially-private data [24] The formal definition is as follows:
Definition 2.3.
Post-Processing Property in ${\epsilon}$-LDP. The post-processing property[23] of LDP states that if a randomized algorithm $\mathcal{A}$ satisfies ${\epsilon}$-LDP, then for any deterministic or randomized function $g$, the composed mechanism $g(\mathcal{A}(\cdot))$ also satisfies ${\epsilon}$-LDP. Specifically, for all $v_{1},v_{2}$ and for all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$, we have:
@@
Post-processing immunity guarantees that applying function $g$ to the output of $\mathcal{A}$ does not compromise the privacy guarantees provided by ${\epsilon}$-LDP.
We prove that LDPKiT’s noise injection scheme satisfies the definition of ${\epsilon}$-LDP with the Laplacian mechanism in Appendix A.

### v3

# LDPKiT: Superimposing Remote Queries for Privacy-Preserving Local Model Training

@@ Abstract
Users of modern Machine Learning (ML) cloud services face a privacy conundrum—on one hand, they may have concerns about sending private data to the service for inference, but on the other hand, for specialized models, there may be no alternative but to use the proprietary model of the ML service. In this work, we present LDPKiT, a framework for non-adversarial, privacy-preserving model extraction that leverages a user’s private in-distribution data while bounding privacy leakage. LDPKiT introduces a novel superimposition technique that generates approximately in-distribution samples, enabling effective knowledge transfer under local differential privacy (LDP). Experiments on Fashion-MNIST, SVHN, and PathMNIST demonstrate that LDPKiT consistently improves utility while maintaining privacy, with benefits that become more pronounced at stronger noise levels. For example, on SVHN, LDPKiT achieves nearly the same inference accuracy at $\epsilon=1.25$ as at $\epsilon=2.0$, yielding stronger privacy guarantees with less than a 2% accuracy reduction. We further conduct sensitivity analyses to examine the effect of dataset size on performance and provide a systematic analysis of latent space representations, offering theoretical insights into the accuracy gains of LDPKiT.
I Introduction
In specialized domains such as medical imaging and cybersecurity, users who wish to employ machine learning (ML) solutions may have no choice but to use an ML model trained by a service provider, such as IBM Watson Health and Darktrace [1, 2, 3]. Because of the lack of publicly available datasets in such privacy-sensitive areas, organizations that have or can acquire access to large proprietary datasets are able to train ML models that average users or the open-source community simply cannot replicate. As a result, when using such a model, a user may have no choice but to send their inputs to the ML service provider for inference, opening themselves to a loss of privacy.
Unfortunately, many ML service providers do not prioritize the privacy of their users. In fact, some service providers explicitly state in their privacy policies that user-provided inference inputs may be stored indefinitely for purposes such as model training and advertising. Additionally, they may share these inputs with third parties to facilitate external services. [4, 5, 6, 7]. Such data practices can directly violate privacy regulations like the European Union’s General Data Protection Regulation (GDPR) [8]. Furthermore, several cases have emerged where providers have misused user data [9, 10, 11, 12]. In 2016, for example, Yahoo secretly colluded with the US government’s digital communication surveillance and used its custom spam and child pornography detection system to monitor user emails [9, 10, 11]. In 2022, the Amazon Ring Doorbell was found to have disclosed users’ video and audio footage to the police without authorization [12]. Even the recent large language models (LLMs) have incurred data leakage [13, 14], leading to many organizations banning their use.
@@ I Introduction
Various methods for protecting privacy during inference have been proposed. For instance, homomorphic encryption [15, 16] and hardware-enforced trusted execution environments [17] ensure that the user’s inputs are always encrypted during inference, so that the service provider cannot see their contents. However, they cannot protect against a malicious or faulty model that leaks information about the inputs. Moreover, they require the cooperation of the ML service provider to deploy such privacy-preserving measures.
The strongest privacy guarantee a user can obtain is to perform model inference locally, so that the user’s private data never leaves their control. However, in domains where training data is scarce and open-source models do not exist, this is not possible. Moreover, the current business model of many ML companies is to deliberately keep their models, code and training data proprietary, so that they may maintain a competitive advantage [18, 19, 20]. They argue that proprietary models support the significant investment they make in improving their models, including acquiring high-quality proprietary datasets for training [21, 22]. As a result, it is also not possible for users to obtain proprietary models from ML service providers to run locally. In such circumstances, users may resort to model extraction: performing queries on the ML service provider’s model and using the results to train their own local model. While ML service providers generally do not condone model extraction, it is difficult for them to stop, and similar techniques, such as result caching to improve edge-device performance, are widely adopted as standard optimizations [23]. Moreover, model extraction still requires users to perform and pay for queries made to the ML service provider, so the provider may tolerate model extraction so long as their business model remains viable.
Motivated by these observations, we propose a model extraction-inspired privacy-preserving technique, LDPKiT (Local Differentially-Private and Utility-Preserving Inference via Knowledge Transfer), which enables users to perform a privacy-protective model extraction on specialized models. However, effective extraction requires users’ queries to be representative of the model’s training data distribution, so ideally, they should use real data. Due to the scarcity of public data in privacy-sensitive domains, users must use their own real data for extraction. Thus, LDPKiT uses local differential privacy (LDP) to bound privacy leakage by injecting random noise into inference queries. Further, users will likely only have a small amount of unlabelled in-distribution data; otherwise, they would be able to train a model themselves. LDPKiT addresses this issue with a novel data augmentation technique that superimposes noised input pairs, generating more in-distribution queries for model extraction. Such a setting naturally arises in practice, for example, in small hospitals with limited but sensitive in-distribution medical scans, or in financial institutions with private in-distribution transaction records. In these scenarios, LDPKiT provides a mechanism for leveraging powerful remote models while ensuring that sensitive user data remains protected.
@@
We acknowledge that model extraction is controversial. In an extreme case, an adversary that extracts an ML service provider’s commercial model with the intention to resell that model and compete with the ML service provider (i.e., model theft) is considered adversarial and may be in direct violation of the provider’s terms of service. We naturally do not advocate such uses of model extraction. In contrast, we focus on privacy protection and employ extraction techniques in a non-adversarial way. Our approach differs from competitive model theft in several important ways: 1) LDPKiT’s extracted model is not competitive with the provider’s model due to the addition of LDP noise and would not violate the terms of service of most commercial model providers [24]. 2) Users perform a large number of queries to the model provider and presumably compensate the model provider for these queries. Since preventing model extraction, especially for small models, is inherently challenging, it is often accounted for in query pricing structures [25, 26, 27].
@@
RQ1. Does LDPKiT recover the utility impacted by LDP noise? (Section IV-B)
RQ2. Why superimposition works in LDPKiT? (Section IV-C)
@@
RQ3. How do the size of the private dataset and the number of queries impact LDPKiT? (Section IV-D)
@@
Contributions. LDPKiT generates inference datasets that are both privacy-protective and close to the distribution of sensitive datasets, allowing users to accurately annotate their private data with bounded privacy leakage. We evaluate LDPKiT across diverse applications and multiple model architectures, systematically examining its performance and sensitivity to dataset sizes. Additionally, we conduct in-depth analyses of latent space representations to uncover factors contributing to LDPKiT’s effectiveness.
II Overview
@@
Figure 1: Comparison of accuracies on SVHN and Fashion-MNIST: Extracting a local model (ResNet-18) with OOD public data samples versus querying the remote model (ResNet-152) with $\epsilon$-LDP protected private data.
II-A Motivation
Our scenario addresses the use of ML services for labelling privacy-sensitive data. We aim to protect the user’s data when querying a service provider’s model during inference, where the untrusted service provider may maliciously process user inputs without consent, violating regulations such as GDPR. To mitigate privacy risks, users without access to sufficient labelled data with which to train their models may opt to extract the service provider’s model. However, successful extraction requires in-distribution data, and relevant public data is usually scarce or unavailable in privacy-sensitive applications. Extracting a model with out-of-distribution (OOD) data fails to produce a functional model. We illustrate this with a simple experiment: we attempt to extract a ResNet-18 model trained on the Fashion-MNIST dataset using samples from SVHN, and, conversely, extract a model trained on the SVHN dataset using samples from Fashion-MNIST. As shown in the two leftmost bars of each group in Figure 1, the resulting accuracies on both datasets are close to random guessing (approximately 10% for these 10-class datasets). In contrast, extracting a model with in-distribution samples, even with noise added to those samples, results in higher accuracy. Specifically, ResNet-18 models trained on noisy in-distribution data, modified using $\epsilon$-LDP with strong privacy guarantees (i.e., low $\epsilon$ values), achieve an average accuracy of around 20%. Although 20% accuracy does not represent a successful model extraction, this result motivates the adoption of $\epsilon$-LDP as the foundation for our privacy-preserving framework. We subsequently enhance this baseline with techniques designed to recover the accuracy degradation incurred from LDP noise. These enhancements yield a fourfold improvement in accuracy under the same privacy level.
II-B Preliminaries and General Setup
There are two parties involved in our scenario: a model provider with the remote model and a user with a small set of unlabelled sensitive data. We use $\mathcal{M}_{\rm R}$}} to denote the remote model and $\mathcal{M}_{\rm L}$}} to denote the model extracted by the user. $\mathcal{D}_{\rm priv}$} denotes the privacy-sensitive dataset the user owns and wishes to label and protect, and $|{\mathcal{D}_{\rm priv}}|$ represents the number of samples in $\mathcal{D}_{\rm priv}$}. To protect $\mathcal{D}_{\rm priv}$}, we add $\epsilon$-LDP noise to each sample, resulting in $\mathcal{D}_{\rm protected}$} (and naturally $|{\mathcal{D}_{\rm priv}}|=|{\mathcal{D}_{\rm protected}}|$). $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm protected}}|$ are generally expected to be small. We thus propose to augment the number of privacy-protected samples available for model extraction, creating an expanded inference dataset, $\mathcal{D}_{\rm infer}$}. The size of $\mathcal{D}_{\rm infer}$} ($|{\mathcal{D}_{\rm infer}}|$) is determined by the user’s query budget, i.e., the number of queries the user is willing to make to $\mathcal{M}_{\rm R}$}, given the pricing model of the service. This allows users to flexibly select a subset or the entire set of augmented samples generated from $\mathcal{D}_{\rm protected}$}. These augmented samples are labelled by querying $\mathcal{M}_{\rm R}$} and subsequently used to train $\mathcal{M}_{\rm L}$}. We use SIDP to denote a baseline use of LDP that does not involve model extraction. In SIDP, the user simply queries $\mathcal{M}_{\rm R}$} with $\mathcal{D}_{\rm protected}$} and accepts the returned results as correct, despite potential errors introduced by the LDP noise in $\mathcal{D}_{\rm protected}$}. The two rightmost bars in Figure 1 show the performance of $\mathcal{M}_{\rm R}$} a user can expect when using SIDP. Currently, LDPKiT supports only the image modality, where superimposition maps naturally to pixel-level operations. Extending this concept to text or other data types would require new mechanisms (e.g., embedding-level or token-level mixing), which we leave as future work. To improve the prediction accuracy on $\mathcal{D}_{\rm priv}$}, LDPKiT proposes to extract an $\mathcal{M}_{\rm L}$} with samples from $\mathcal{D}_{\rm infer}$} generated using two data augmentation methods: LDPKiT-Rand and LDPKiT-Sup, which we discuss in Section III-B.
II-C Threat Model
The user’s goal is to label samples in $\mathcal{D}_{\rm priv}$} with reasonable accuracy when interacting with an ML service in a privacy-preserving way. We assume that $\mathcal{M}_{\rm R}$} is honest but curious. It honestly answers the user’s queries but may record the queries and results to infer information about the user. To mitigate this threat, the user applies $\epsilon$-LDP to their data before querying $\mathcal{M}_{\rm R}$}, limiting the amount of information that can be inferred from individual samples. The user knows $\mathcal{M}_{\rm R}$}’s task and functionality, but has no access to its underlying model architecture and dataset on which it was trained. $\mathcal{M}_{\rm L}$} aims to learn $\mathcal{M}_{\rm R}$}’s functionality, relying only on the black-box access limited to hard label responses from query interactions.
II-D Privacy Guarantee
@@
LDPKiT ensures the privacy of $\mathcal{D}_{\rm priv}$} during inference by applying LDP noise to each data sample. This section discusses the key concepts and methodologies essential for designing effective privacy-preserving data augmentation mechanisms.
Definition II.1.
@@
A lower $\epsilon$ value indicates a tighter bound of the equation and a stronger privacy guarantee. We choose $\epsilon$ values primarily based on the industry standards [29, 30] and empirically verify that privacy-protected data is resilient to data reconstruction attacks [31]. We apply Laplacian noise to satisfy the $\epsilon$-LDP privacy guarantee, and the definition is as follows.
Definition II.2.
@@
For the design of the candidate dataset used for inference, $\mathcal{D}_{\rm infer}$}, we also leverage LDP’s post-processing property to facilitate utility recovery. If an algorithm safeguards an individual’s privacy, a data analyst cannot compromise privacy—whether under the formal definition or in any intuitive sense—merely by reflecting on the algorithm’s output [32]. Post-processing is commonly used in DP schemes to improve the interpretability or accuracy of differentially-private data analysis [33]. The formal definition is as follows.
Definition II.3.
Post-Processing Property in $\epsilon$}-LDP. The post-processing property of LDP [32] states that if a randomized algorithm $\mathcal{A}$ satisfies $\epsilon$}-LDP, then for any deterministic or randomized function $g$ that does not depend on the original input values, the composed mechanism $g(\mathcal{A}(\cdot))$ also satisfies $\epsilon$}-LDP. Specifically, for all $v_{1},v_{2}$ (i.e., all values in the domain) and for all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$:
@@
The post-processing property ensures that applying a dataset-independent function $g$ to the output of $\mathcal{A}$ cannot weaken the privacy guarantees provided by $\epsilon$}-LDP, as long as $g$ itself does not access the raw data.